# CANVAS: Commonsense-Aware Navigation System for Intuitive Human-Robot Interaction

### v2

@@ I Introduction
In response to these challenges, we introduce CANVAS (Commonsense-Aware NaVigAtion System), a novel framework for integrating abstract human instructions into robot navigation. Our approach utilizes both visual and linguistic inputs, such as rough sketch trajectories on map images or textual descriptions. These multimodal instructions are processed by a vision-language model that generates incremental navigation targets. By leveraging the commonsense knowledge embedded in pre-trained vision-language models [11, 12, 13, 14], robots can develop a versatile understanding of commonsense navigation dynamics. Quantifying successful navigation behaviors using rewards[15, 16, 17] is particularly difficult in commonsense-aware navigation. Therefore, we employ imitation learning, enabling the robot to comprehend user intent behind noisy and imprecise instructions from human demonstrations. [18]
@@ III-B Task Definition
At each timestep $t$, the robot $R$ manages two states: the front view image $X_{f}(t)$ and the robot’s hindsight trajectory [44] up to timestep $t-1$, denoted as $H(t)$. The front view image $X_{f}(t)$ is captured by the robot’s camera, while $H(t)$ is tracked through odometry to log the robot’s past positions. We combine the sketch instruction $S$ and hindsight trajectory $H(t)$ onto the same map to create the canvas map image $X_{c}(t)$. At each step, the robot $R$ generates an action $y(t)=[w_{0},w_{1},w_{2},w_{3}]$, which is a sequence of waypoints. This action is conditioned on the front view image $X_{f}(t)$, the canvas map image $X_{c}(t)$, and the language instruction $L$. Formally, the robot’s action is defined as:
@@
At the end of each iteration, the hindsight trajectory is updated by appending $p(t)$, which represents the robot’s position updated through predicted waypoints $y(t)$, resulting in $H(t+1)=(p(1),p(2),...,p(t))$. This update is reflected in the next canvas map image $X_{c}(t+1)$, while the front view image $X_{f}(t+1)$ is also updated based on the robot’s new position. This process continues until the robot either reaches the destination or a maximum timestep $t=T$ is reached.
@@ IV-A Architecture
The model architecture is illustrated in Figure 3. We adopt a VLM denoted as $\pi_{\theta}$. The front view image $X_{f}(t)$ and canvas map image $X_{c}(t)$ are processed through a vision encoder $g_{\phi}(\cdot)$. This results in two visual features, $Z_{f}=g_{\phi}(X_{f}(t))$ and $Z_{c}=g_{\phi}(X_{c}(t))$. A projector $p_{\phi}$ is used to project these visual features into the word embedding space, producing a sequence of visual tokens $\tau_{v}=p_{\phi}(Z_{f},Z_{c})$. A sequence of language tokens $\tau_{l}$ is also obtained from the language instruction $L$. Both the visual tokens $\tau_{v}$ and language tokens $\tau_{l}$ are then fed into the large language model denoted as $f_{\phi}(\cdot)$, which outputs the waypoint tokens $[w_{0},w_{1},w_{2},w_{3}]=f_{\phi}(\tau_{v},\tau_{l})$. Due to the reasons mentioned in Section II-B, we apply the simplest method, K-means clustering, to discretize continuous waypoints into tokens, with empirical testing showing that K=128 outperformed 32, 64, or 256. Fewer tokens can hinder precise actions like navigating narrow passages.
@@ V Experiments
We compare CANVAS with ROS NavStack [20], a straightforward yet effective rule-based navigation system. For this system, we converted the sketch instructions into step-by-step, point-to-point inputs, but language instructions could not be accommodated. The same hyperparameters were used for all experiments with NavStack. We evaluate two variations of CANVAS. CANVAS-S modifies the original Idefics2 8B [45] by swapping the vision encoder for SigLIP-L [46] and the text encoder for Qwen2-0.5B [47], reducing the model size from 8B to 0.7B to better accommodate real-world deployment. In contrast, CANVAS-L retains the original Idefics2 8B [45] architecture with its pre-trained weights. Both models were trained using LoRA [48] with $r$=256, $a$=512, and a dropout rate of 0.1. We use the AdamW optimizer [49] with a different learning rates of 2e-5 for the LLM and 5e-5 for the projector and vision encoder, a batch size of 32, and 5 training epochs. Each model utilizes 128 waypoint tokens. In the main experiments, CANVAS models were inferred using a single NVIDIA H100 GPU. All experiments were evaluated over three iterations for each test dataset, with randomized starting orientations. In the real-world environment, SLAM with FAST-LIO2 [50] was used to find the robot’s current position.
@@ V-A Results in the Simulated Environments
Seen Environments. Table V shows CANVAS’s performance in three seen environments: office, street (road, sidewalk), and orchard. Under precise instructions, CANVAS achieves similar SR and CR to NavStack in the office and street (road), where navigation is easier, indicating that CANVAS can learn the essential navigation behaviors effectively from human demonstrations. However, in more challenging environments like the street (sidewalk) and orchard, CANVAS significantly outperforms NavStack. A detailed analysis of CANVAS’s performance is provided in Section V-D.
@@
Unseen Environment. We exclude the gallery environment during training to evaluate CANVAS’s performance in unseen settings. As demonstrated in Table V, CANVAS continues to show strong navigation capabilities, even in scenarios with noisy guidance.

### v3

@@ I Introduction
In response to these challenges, we introduce CANVAS (Commonsense-Aware NaVigAtion System), a novel framework for integrating abstract human instructions into robot navigation. Our approach utilizes both visual and linguistic inputs, such as rough sketch trajectories on map images or textual descriptions. These multimodal instructions are processed by a vision-language model that generates incremental navigation targets. By leveraging the commonsense knowledge embedded in pre-trained vision-language models [11, 12, 13, 14, 15], robots can develop a versatile understanding of commonsense navigation dynamics. Quantifying successful navigation behaviors using rewards[16, 17, 18] is particularly difficult in commonsense-aware navigation. Therefore, we employ imitation learning, enabling the robot to comprehend user intent behind noisy and imprecise instructions from human demonstrations. [19]
@@ III-B Task Definition
At each timestep $t$, the robot $R$ manages two states: the front view image $X_{f}(t)$ and the robot’s hindsight trajectory [45] up to timestep t1t-1, denoted as $H(t)$. The front view image $X_{f}(t)$ is captured by the robot’s camera, while $H(t)$ is tracked through odometry to log the robot’s past positions. We combine the sketch instruction $S$ and hindsight trajectory $H(t)$ onto the same map to create the canvas map image $X_{c}(t)$. At each step, the robot $R$ generates an action $y(t)=[w_{0},w_{1},w_{2},w_{3}]$, which is a sequence of waypoints. This action is conditioned on the front view image $X_{f}(t)$, the canvas map image $X_{c}(t)$, and the language instruction $L$. Formally, the robot’s action is defined as:
@@
At the end of each iteration, the hindsight trajectory is updated by appending $p(t)$, which represents the robot’s position updated through predicted waypoints $y(t)$, resulting in H(t+1)=(p(1),p(2),,p(t))H(t+1)=(p(1),p(2),...,p(t)). This update is reflected in the next canvas map image $X_{c}(t+1)$, while the front view image $X_{f}(t+1)$ is also updated based on the robot’s new position. This process continues until the robot either reaches the destination or a maximum timestep $t=T$ is reached.
@@ IV-A Architecture
The model architecture is illustrated in Figure 3. We adopt a VLM denoted as $\pi_{\theta}$. The front view image $X_{f}(t)$ and canvas map image $X_{c}(t)$ are processed through a vision encoder $g_{\phi}(\cdot)$. This results in two visual features, $Z_{f}=g_{\phi}(X_{f}(t))$ and $Z_{c}=g_{\phi}(X_{c}(t))$. A projector $p_{\phi}$ is used to project these visual features into the word embedding space, producing a sequence of visual tokens $\tau_{v}=p_{\phi}(Z_{f},Z_{c})$. A sequence of language tokens $\tau_{l}$ is also obtained from the language instruction $L$. Both the visual tokens $\tau_{v}$ and language tokens $\tau_{l}$ are then fed into the large language model denoted as $f_{\phi}(\cdot)$, which outputs the waypoint tokens $w_{0},w_{1},w_{2},w_{3}]=f_{\phi}(\tau_{v},\tau_{l})$. Due to the reasons mentioned in Section II-B, we apply the simplest method, K-means clustering, to discretize continuous waypoints into tokens, with empirical testing showing that K=128 outperformed 32, 64, or 256. Fewer tokens can hinder precise actions like navigating narrow passages.
@@ V Experiments
We compare CANVAS with ROS NavStack [21], a straightforward yet effective rule-based navigation system. For this system, we converted the sketch instructions into step-by-step, point-to-point inputs, but language instructions could not be accommodated. The same hyperparameters were used for all experiments with NavStack. We evaluate two variations of CANVAS. CANVAS-S modifies the original Idefics2 8B [46] by swapping the vision encoder for SigLIP-L [47] and the text encoder for Qwen2-0.5B [48], reducing the model size from 8B to 0.7B to better accommodate real-world deployment. In contrast, CANVAS-L retains the original Idefics2 8B [46] architecture with its pre-trained weights. Both models were trained using LoRA [49] with rr=256, aa=512, and a dropout rate of 0.1. We use the AdamW optimizer [50] with a different learning rates of 2e-5 for the LLM and 5e-5 for the projector and vision encoder, a batch size of 32, and 5 training epochs. Each model utilizes 128 waypoint tokens. In the main experiments, CANVAS models were inferred using a single NVIDIA H100 GPU. All experiments were evaluated over three iterations for each test dataset, with randomized starting orientations. In the real-world environment, SLAM with FAST-LIO2 [51] was used to find the robot’s current position.
@@ V-A Results in the Simulated Environments
Seen Environments. Table III shows CANVAS’s performance in three seen environments: office, street (road, sidewalk), and orchard. Under precise instructions, CANVAS achieves similar SR and CR to NavStack in the office and street (road), where navigation is easier, indicating that CANVAS can learn the essential navigation behaviors effectively from human demonstrations. However, in more challenging environments like the street (sidewalk) and orchard, CANVAS significantly outperforms NavStack. A detailed analysis of CANVAS’s performance is provided in Section V-D.
@@
Unseen Environment. We exclude the gallery environment during training to evaluate CANVAS’s performance in unseen settings. As demonstrated in Table III, CANVAS continues to show strong navigation capabilities, even in scenarios with noisy guidance.